If the following reaction was used to produce hydrogen gas and the carbon dioxide is vented to the atmosphere, what is the % atom economy? CH4+2 H2O--->CO2+4 H2
Answer #1

CH4+2 H2O--->CO2+4 H2

Using the atomic masses of H = 1, C = 12, O = 16, we can calculate the atom economy .

the reaction equation can be expressed in terms of theoretical reacting mass units

[(12 ) + 4(1)] + [2 x (2 + 16)] ===> [12 +2 x 16] + [4 x (2)]

[16 of CH4] + [36 of H2O] ===> [44 of CO2] + [8 of H2]

so there are a total of 8 mass units of the useful/desired product H2

out of a total mass of reactants or products of 16 + 36 = 44 + 8 = 104

   TOTAL FORMULA MASSES of USEFUL PRODUCT

ATOM ECONOMY   = 100 x  

-----------------------------------------------------------------------------------
      TOTAL FORMULA MASSES of all REACTANTS

Therefore the atom economy = 100 x 8 / 104 = 7.69 %

I JUST COPIED FORMULA... HOPE ITS CORRECT